radomes.openskiesproject.org
ďfrrdrrtf5c9 ďfrrdrrtf5c9
  • 10-03-2017
  • Mathematics
Answered

a game has three possible outcomes : win, lose, or tie. The probability of a loss is 0.1, and the probability of the tie is 0.5. on a single play of the game the probability of NOT winning is

a) 0.1.
b) 0.33.
c) 0.5.
d) 0.6.

Verified Answer :

Anonymous Anonymous
  • 12-03-2017

P(not winning) = p(loss) + P(tie) = 0.1 + 0.5 = 0.6

Answer Link

Related Questions

Explain the scope of own role in supporting an individual to negotiate an environment
Who was the founder of the religion Buddhism
What is the chain of events driven by conflict and decision making in a text plot characterization setting theme
What is 15 of 34 Please explain
A solution has a OH of 1 x 1012 What is the pOH of this solution 2 7 10 12
What is a way that you can write 83041 in notation form
how did the romans keep their large empire together
Why was the term political machine used to describe corrupt political systems in the late 1800s Political machines ran very smoothly like clocks Political machi
Which statement is correct A All organisms that can synthesize their food by means of photosynthesis are heterotrophic B All organisms that cannot synthesize th
What is the molar mass of ph